# Class-Conditional Distribution Balancing for Group Robust Classification

### v2

@@ II Related Work
II-A Bias-guided Methods
@@
Addressing spurious correlations becomes straightforward when bias annotations are available. By combining both bias and label annotations, one can create a group balanced training process or model selection strategy, to enhance robustness against spurious correlations. For example, Group distributionally robust optimization (GroupDRO) [22, 43, 44, 45] proposes to minimize the worst-group loss based on predefined group assignments. To alleviating the reliance on bias annotations, deep feature reweighting (DFR) [27, 46] proposes retraining only the last layer of the model using a small, group-balanced held-out dataset. MAPLE [28] utilizes a validation set with bias annotations to reweight the training samples, ensuring that the training process focuses on the worst group. Pseudo-bias-supervised methods propose to use the predicted bias information instead of explicit annotations, typically through auxiliary models. Learning from Failure (LFF) [24], logit adjustment technique (LC) [26] and Disentangled Feature Augmentation (DFA) [47] reweight the loss of the unbiased model using a co-trained biased model. Meanwhile, Just-train-twice (JTT) [25] reweights misclassified samples from an initial biased training phase, emphasizing data from the worst-performing groups. Additionally, research in [23, 48, 49] integrates large-scale vision-language models for more precise bias prediction. However, the performance of these methods depends on the prediction quality and still requires bias annotations during validation for optimal performance.
II-B Bias-free Methods
To eliminate the dependence on bias annotations/predictions, researchers have developed bias-free methods under different motivations. Stable learning based methods [50] propose to decorrelate all features to exclude spurious correlations. For example, [42, 51] proposes a sample reweighting approach to decorrelate input variables and demonstrate stable performance across distribution shifts. StableNet [5] further generalizes these techniques to nonlinear feature decorrelation. However, achieving perfect decorrelation theoretically requires a sufficiently large dataset, which is often impractical. To compensate for the imperfectness of sample reweighting under finite-sample settings, Sparse variable independent(SVI) [40] introduces a sparsity constraint to avoid unnecessary decorrelation within causal variables. Most of these methods are developed within linear frameworks with predefined features, and their extension to deep models is less effective. This is because the learned features from deep models are often heavily entangled, making decorrelation insufficient to disentangle causal and spurious factors [40, 52]. Another line of research utilizes disagreements among diverse models to improve generalization under distribution shifts. For instance, Diversity by disagreement [53] and Diversify and disambiguate [54] propose to maximize the disagreement between multiple predictors to learn a diverse ensemble, which require a completely spuriously correlated training dataset. Selective last-layer finetuning (SELF) [2] proposes reweighting the dataset based on misclassifications or disagreements observed before and after last-layer retraining. However, these methods still need a small set of group annotations for model selection. Several other bias-free methods have demonstrated competitive performance, e.g., General Greedy De-bias Learning [55] leverages both a biased model and a base model to assess the difficulty of training samples, thereby focusing training on the more challenging ones. However, this approach may overestimate the level of bias, necessitating the use of curriculum regularization to strike a better balance between in-distribution and out-of-distribution performance. uLA [8] employs biased prediction to adjust the logits for debias learning. It relies heavily on extensively pretrained self-supervised models, which significantly influence the final performance. XRM [29] constructs twin networks to fully predict bias across the entire training dataset, using a label-flipping mechanism to intentionally steer the model toward biased predictions. MASKTUNE [30], on the other hand, explores new features during a secondary training phase by masking previously identified spurious factors. However, it requires careful selection of the masking threshold to ensure stable performance. The work presented in [56] introduces a novel perspective on the causes of spurious correlations. While it shares certain similarities with our approach, it is grounded in a different framework and primarily aims to offer a comprehensive analysis of existing methods.
Our work shares similarities with stable learning in its bias-free approach and the use of sample reweighting techniques to adjust the distribution of training data. The key differences lie in: ($i$)our modified reweighting strategy enables balancing at both the distribution and class levels; ($ii$) instead of decorrelating all features, we take a more straightforward approach by minimizing the mutual information between spurious factors and class labels; ($iii$) our method is fully compatible with deep learning and has no limitations.
III Our Method
Given each observation consists of an input $\boldsymbol{x}$ (e.g., an image $\boldsymbol{x}\in\mathbb{R}^{K\times H\times W}$) and a label $y\in\mathbb{R}$, where $K$, $H$, and $W$ represent the channel, height and width of the image, respectively, the spurious correlation problem is defined as follows. Let $\boldsymbol{z}\in\mathbb{R}^{D}$ denote the latent representation of $\boldsymbol{x}$ after feature extraction, which contains causal and spurious factors, $\boldsymbol{z}=[\boldsymbol{z}_{\text{c}},\boldsymbol{z}_{\text{s}}]$. Here, $\boldsymbol{z}_{\text{c}}\in\mathbb{R}^{D_{\text{c}}}$ and $\boldsymbol{z}_{\text{s}}\in\mathbb{R}^{D_{\text{s}}}$ are implicitly entangle and technically challenging to decorrelate. $D$, $D_{\text{c}}$, and $D_{\text{s}}$ denote the dimensionalities of the corresponding latent variables. Theoretically, $\boldsymbol{z}_{\text{c}}$ represents the true causal factors that define the class labels and thus strongly correlated with $y$, whereas $\boldsymbol{z}_{\text{s}}$ consists of spurious factors unrelated to class information. However, under distribution shift, the presence of unobservable confounders [57, 58] induces a strong spurious correlation between $\boldsymbol{z}_{\text{s}}$ and $y$ in the training data [59, 39], which fails to hold during test. Such unintended spurious correlations undermine model generalization. To address this issue, we propose breaking spurious correlations from the view of the joint distribution $p(\boldsymbol{z}_{\text{s}},y)$ and providing a debiased data distribution for classification by directly minimizing the mutual information $I(\boldsymbol{z}_{\text{s}},y)$ as follows,
$I(\boldsymbol{z}_{\text{s}},y)$ $=\mathbb{E}_{p(y)}\mathrm{KL}[p(\boldsymbol{z}_{\text{s}}|y)\|p(\boldsymbol{z}_{\text{s}})]$ (1)
$=\mathbb{E}_{p(y,\boldsymbol{z}_{\text{s}})}\log\frac{p(y|\boldsymbol{z}_{\text{s}})}{p(y)}$
$=H(y)-H(y|\boldsymbol{z}_{\text{s}})$
which considers both linear and non-linear dependency between $\boldsymbol{z}_{\text{s}}$ and $y$. According to the fact that: ($i$) $H(y)\geq H(y|\boldsymbol{z}_{\text{s}})$; ($ii$) $y$ is discrete, and the entropy of a discrete random variable is non-negative [60]; we have the following conclusion, minimizing the mutual information $I(\boldsymbol{z}_{\text{s}},y)$ is equivalent to maximize $H(y|\boldsymbol{z}_{\text{s}})$, namely, minimize $-H(y|\boldsymbol{z}_{\text{s}})$, detailed as follows:
$\mathcal{L}_{\text{mutual}}$ $=-H(y|\boldsymbol{z}_{\text{s}})$ (2)
$=I(\boldsymbol{z}_{\text{s}},y)-H(y)$
Intuitively, Equation 2 shows that spurious correlations arise from two primary factors: ($i$) the mismatch among the class-conditional distributions, termed distribution imbalance and ($ii$) the non-uniformity of the class distribution, termed class imbalance. Motivated by these key insights, we propose the Class-conditional distribution balancing (CCDB) technique to break spurious correlations by minimizing Equation 2, which involves reducing the distributional distance between $p(\boldsymbol{z}_{\text{s}}|y)$ and $p(\boldsymbol{z})$ while maximizing the entropy of $p(y)$. Note, the first term in Equation 2 drives the distribution of $\boldsymbol{z}_{\text{s}}$ within each class toward a common marginal distribution, perceptually a balancing at the distributional level. The second term promotes uniformity across classes, ensuring class balance. Thus, Equation 2 generalizes traditional class-balancing methods by addressing both distributional and class-level imbalances within a unified framework.
Directly minimizing Equation 2 is nontrivial due to the unavailability of the disentangled $\boldsymbol{z}_{\text{s}}$ and its exact distributions. For practical implementation, we first propose learning a biased feature extractor in Section III-B, which captures features dominated by spurious correlations. This enables us to replace $\boldsymbol{z}_{\text{s}}$ with $\boldsymbol{z}_{\text{c}}$, as we do in the following sections. Secondly, we introduce the following designations to handle each term in Equation 2 separately.
$W_{2}(\boldsymbol{z}|_{y=c},\boldsymbol{z})=\|\boldsymbol{\mu}_{c}-\overline{\boldsymbol{\mu}}\|_{2}^{2}+d({\bf C}_{c},\overline{{\bf C}})\$ (3)
where, $d({\bf C}_{c},\overline{{\bf C}})=\text{trace}({\bf C}_{c}+\overline{{\bf C}}-2[\overline{{\bf C}}^{\frac{1}{2}}{\bf C}_{c}\overline{{\bf C}}^{\frac{1}{2}}]^{\frac{1}{2}})$, $\{\boldsymbol{\mu}_{c},{\bf C}_{c}\}$ and $\{\overline{\boldsymbol{\mu}},\overline{{\bf C}}\}$ denote the mean and covariance matrix for $p(\boldsymbol{z}|y=c)$ and $p(\boldsymbol{z})$, respectively.
The entropy term. This term can be easily maximized through data subsampling or reweighting, following traditional class balancing techniques [63]. In fact, it is possible to handle these two terms together. In the following section, we propose a modified sample reweighting mechanism to achieve this goal.
III-A Intra-class Sample Reweighting for CCDB
Figure 1: Analyzing the effectiveness of our method under three extreme configurations where $\boldsymbol{z}_{\text{s}}$ and $y$ are: (a) strongly correlated with well-disentangled predefined features; (b) completely independent; and (c) strongly correlated with predefined features that are independent but semantically entangled. Each case consists of two classes (marked in blue and red). Data points in the yellow regions are assigned with high weights, while those in the blue regions are assigned with low weights.
Inspired by the success of sample reweighting in decorrelation [5, 40], we propose leveraging it to remove the dependence between $\boldsymbol{z}$ and $y$. Unlike existing methods that learn a reweighting scheme over the entire training data, our approach applies sample reweighting independently within each class, which enables automatic optimization of both terms in Equation 2.
Suppose the observations are from $C$ classes, with each class represented as $\{\boldsymbol{x}_{i},y=c\}_{i=1}^{n_{c}}$, where $n_{c}$ denotes the sample count for class $c$. We introduce class-specific sample weights to adjust the class-conditional distribution. Specifically, for class $c$, we define a weighting vector $\boldsymbol{w}_{c}\in\mathbb{R}^{n_{c}}$, satisfying $[\boldsymbol{w}_{c}]_{i}\geq 0$ for $i=1,\cdots,n_{c}$, and $\sum_{i=1}^{n_{c}}[\boldsymbol{w}_{c}]_{i}=1$. After reweighting, the distribution distance between $p(z|y=c,\boldsymbol{w}_{c})$ and $p(z)$ following Equation 3 is computed as:
$W_{2}(\boldsymbol{z}|_{\{y=c,\boldsymbol{w}_{c}\}},\boldsymbol{z})=\|\boldsymbol{\mu}_{c}(\boldsymbol{w}_{c})-\overline{\boldsymbol{\mu}}\|_{2}^{2}+d({\bf C}_{c}(\boldsymbol{w}_{c}),\overline{{\bf C}})$ (4)
where $\boldsymbol{\mu}_{c}(\boldsymbol{w}_{c})$ and ${\bf C}_{c}(\boldsymbol{w}_{c})$ denote the mean and covariance of class $c$ after reweighting, defined as,
$\boldsymbol{\mu}_{c}(\boldsymbol{w}_{c})$ $=\sum_{i=1,\boldsymbol{y}_{i}=c}^{n_{c}}[\boldsymbol{w}_{c}]_{i}\cdot\boldsymbol{z}_{i}$ (5)
${\bf C}_{c}(\boldsymbol{w}_{c})$ $=\sum_{i=1,\boldsymbol{y}_{i}=c}^{n_{c}}[\boldsymbol{w}_{c}]_{i}\cdot(\boldsymbol{z}_{i}-\boldsymbol{\mu}_{c})^{T}(\boldsymbol{z}_{i}-\boldsymbol{\mu}_{c})$
Note that we choose to set $\overline{\boldsymbol{\mu}}$ and $\overline{{\bf C}}$ as constants rather than functions of $\{\boldsymbol{w}_{1},\cdots,\boldsymbol{w}_{C}\}$ for the following two reasons: ($i$) The marginal distribution $(\overline{\boldsymbol{\mu}},\overline{{\bf C}})$ serves as an anchor to stably guide the class-conditional distributions toward it. In the ideal case where all class-conditional distributions match the marginal distribution, the marginal distribution will remain unchanged. ($ii$) If, instead, $\overline{\boldsymbol{\mu}}$ and $\overline{{\bf C}}$ were functions of $\{\boldsymbol{w}_{1},\cdots,\boldsymbol{w}_{C}\}$, the objective would become overly flexible and could lead to degenerate solutions—for example, allowing for many possible marginal distributions.
Accordingly, the objective in Equation 2 is reformulated as follows,
$\mathcal{L}_{\text{mutual}}({\boldsymbol{w}_{1},\cdots,\boldsymbol{w}_{C}})=\frac{1}{C}\sum_{c=1}^{C}W_{2}(\boldsymbol{z}|_{\{y=c,\boldsymbol{w}_{c}\}},\boldsymbol{z})$ (6)
where we omit the entropy term because $\boldsymbol{w}_{c}$, being class-specific and summing to $1$, automatically assigns higher weights to classes with fewer samples, thus balancing the long-tailed class distribution. The final objective for our CCDB is:
$\mathop{\min}_{\boldsymbol{w}_{1},\cdots,\boldsymbol{w}_{C}}$ $\frac{1}{C}\sum_{c=1}^{C}W_{2}(\boldsymbol{z}|_{\{y=c,\boldsymbol{w}_{c}\}},\boldsymbol{z})$ (7)
s.t. $\forall c\in\{1,\cdots,C\},\sum_{i=1}^{n_{c}}[\boldsymbol{w}_{c}]_{i}=1,$
$\forall i\in\{1,\cdots,n_{c}\},[\boldsymbol{w}_{c}]_{i}\geq 0,$
The optimal solution of Equation 7 reweights samples within each class by emphasizing those scarcely distributed in areas where the other classes are densely clustered, while understating samples in the opposite scenario. This adjustment aligns the class-conditional distribution to the marginal distribution as closely as possible. In the ideal case, $p(\boldsymbol{z}|y,\boldsymbol{w}_{c})=p(\boldsymbol{z})$, indicating independence between $\boldsymbol{z}$ and $y$. Training a classifier on this reweighted dataset effectively removes reliance on spurious factors.
A toy example is provided in Figure 1 to demonstrate the effectiveness of our CCDB in addressing three extreme situations where the spurious factors and class labels are either strongly dependent(Figure 1(a)(c)) or completely independent (Figure 1(b)). In Figure 1(a) and (c), $90\%$ of the samples have $\boldsymbol{z}_{\text{s}}$ correlated with class labels, while the remaining $10\%$ display conflicting correlations. For each configuration, we draw the following conclusions: (a) Since the data distribution strongly correlates $\boldsymbol{z}_{\text{s}}$ with $y$, ERM places the decision boundary near the minority groups, leading to poor generalization in that region. In contrast, our CCDB (and stable learning) emphasizes the minority area, resulting in a new distribution with the dependency between $\boldsymbol{z}_{\text{s}}$ and $y$ significantly reduced. Thus the decision boundary is determined solely by $\boldsymbol{z}_{\text{c}}$. (b) When $\boldsymbol{z}_{\text{s}}$ is independent of $y$, both ERM and CCDB correctly placed the decision boundary. Additionally, CCDB further emphasizes samples near the boundary and improves performance in ambiguous regions. (c) When predefined features are independent but significantly entangled(e.g., coincidentally aligned with the PCA direction with each derection as a mixture of $\boldsymbol{z}_{\text{c}}$ and $\boldsymbol{z}_{\text{s}}$), independent-based stable learning methods fail and place the decision boundary near the minority area. In contrast, CCDB can still identify and emphasize the minority groups, properly rectifying the decision boundary.
@@ III-A Intra-class Sample Reweighting for CCDB
TABLE I: An illustration of the datasets used in this paper. In the training data, $\boldsymbol{z}_{\text{s}}$ is highly correlated with the class labels. However, this correlation is either broken or even reversed in the test set.
@@
III-B Full Pipeline of Our Method
Stage 1: Learning biased feature extractor.
In this stage, we train a classifier $f_{\boldsymbol{\theta}_{2}}\circ f_{\boldsymbol{\theta}_{1}}$ on $\mathcal{D}_{tr}$, where $f_{\boldsymbol{\theta}_{1}}$ is the backbone and $f_{\boldsymbol{\theta}_{2}}$ is a fully connected classification head. After training, $f_{\boldsymbol{\theta}_{1}}$ is used as a biased feature extractor in the next stage. Intuitively, we aim for the extracted features to be as dominated by spurious factor as possible, allowing us to effectively reduce their impact in the subsequent sample reweighting stage. To achieve this, we introduce the following two strategies:
Data splitting. We experimentally find that the size of the training data significantly affects the dominance of spurious factors. As pointed out in [29], when training on the entire training dataset, the model tends to overfit, causing the latent distribution to be less influenced by biased information. This, in turn, makes our method less effective. Figure 7 shows the latent space learned from different splits of the training data. As the proportion of training data, $\gamma$, decreases, the extracted features become increasingly dominated by spurious factors. Based on this observation, we propose splitting the training data into a subset with proportion $\gamma$ (e.g., $\gamma=50\%$) and training $f_{\boldsymbol{\theta}_{1}}$ on it. The resulting $f_{\boldsymbol{\theta}_{1}}$ is then applied to the entire training set for feature extraction and sample reweighting. Our experiments show that data splitting is critical for fully exposing spurious factors, allowing them to be effectively balanced during the reweighting process.
Intra-class compactness. In addition to classification, we consider to enhance intra-class compactness in the latent space by minimizing the distance between data samples belonging to the same class. The objective for feature extraction is thus reformed as follows:
$+\lambda\frac{1}{C}\sum_{c=1}^{C}\frac{1}{n_{c}}\sum_{i\neq j,y_{i}=y_{j}=c}\|f_{\boldsymbol{\theta}_{1}}(\boldsymbol{x}_{i})-f_{\boldsymbol{\theta}_{1}}(\boldsymbol{x}_{j})\|^{2}$
[…]
Algorithm 1 Class-Conditional Distribution Balancing (CCDB) for group robust classification
  Input: Randomly initialized network $f_{\boldsymbol{\theta}_{2}}\circ f_{\boldsymbol{\theta}_{1}}$ for biased feature extraction, $f_{\boldsymbol{\phi}}$ for classification, logit vectors $\{\boldsymbol{s}_{c}\}_{c=1}^{C}$ for sample reweighting, (initialized with all elements as $-1$, training set $\mathcal{D}_{tr}$ and validation set $\mathcal{D}_{v}$.
  Output: unbiased classifier $f_{\boldsymbol{\phi}}$:
  ##\##Stage1: learning biased feature extractor
  1: Randomly split a subset from $\mathcal{D}_{tr}$ with proportion $\gamma$.
  2: Train $f_{\boldsymbol{\theta}_{2}}\circ f_{\boldsymbol{\theta}_{1}}$ on the split subset using constrained ERM (Equation 8).
  ##\## Stage2: Learning sample weights
  4: Obtain the optimal sample weights $\boldsymbol{w}^{*}=\{\boldsymbol{w}_{1}^{*},\cdots,\boldsymbol{w}_{C}^{*}\}$ by solving Equation 7.
  ##\## Stage 3: Learning unbiased classifier
Stage 2: Learning sample weights.
By optimizing Equation 7, we obtain weights for each training sample that significantly reduce the dependence between spurious factors and class labels. For practical implementation, we reparameterize the weight vector $\boldsymbol{w}_{c}$ for class $c$ as the softmax of a logits vector $\boldsymbol{s}_{c}\in\mathbb{R}^{n_{c}}$, i.e., $\boldsymbol{w}_{c}=\text{softmax}(\boldsymbol{s}_{c})$. This ensures the weights are non-negative and satisfy $\sum_{i=1}^{n_{c}}[\boldsymbol{w}_{c}]_{i}=1$, with the elements of $\boldsymbol{s}_{c}$ being real-valued. To ensure stable optimization, we clip the values of $\boldsymbol{s}_{c}$ within the range $[-T,T]$, where the threshold $T=2$ typically works well.
Stage 3: Learning unbiased classifier.
There are two ways to incorporate sample weights into ERM for unbiased learning. ($i$) Applying the weights during the data sampling process with “torch.utils.data.WeightedRandomSampler”. ($ii$) Applying the sample weights to the cross-entropy loss. We prefer the former, as it empirically yields better performance.
IV Experimental Results
Dataset. We conduct experiments on five widely used benchmark datasets: cMNIST [64, 65], cCIFAR10 [66], Waterbirds [28, 27], CelebA [22, 28], and CivilComments [67] to demonstrate the effectiveness of our method. Among them, cMNIST and cCIFAR10 are ten-way classification tasks, where each class is dominated by a specific color or noise pattern, making them ideal for demonstrating the superior performance of CCDB in challenging multi-class settings with extremely biased training data. Waterbirds, CelebA, and CivilComments are real-world datasets: Waterbirds and CelebA are image classification datasets where each class is strongly correlated with a specific backgrounds or gender attribute; CivilComments is a text classification dataset with the class information spuriously correlated with mentions of certain demographic identities. These datasets are used to showcase the generalizability of CCDB to real-world scenarios. We follow the challenging setup of [25], where no bias/group annotations are available during training or model selection. Details of the data settings are summarized in Table I.
TABLE II: The hyperparameters and augmentations used during the training of our CCDB.
@@ IV Experimental Results
For all datasets, we use the standard train-validation-test splits as defined in their respective literature and report both group-balanced test accuracy and the worst-group test accuracy.
TABLE III: Results on cMNIST and cCIFAR10 with various bias-conflicting ratios in the training set. The test accuracy(%) is averaged over 5 random seeds.The best results are indicated in bold.
Methods Bias label cMNIST cCIFAR10
Train Val $0.5\%$ $1\%$ $2\%$ $5\%$ $0.5\%$ $1\%$ $2\%$ $5\%$
GroupDRO Yes Yes 63.12 68.78 76.30 84.20 33.44 38.30 45.81 57.32
LFF No Yes 52.50$\pm$2.43 61.89$\pm$4.97 71.03$\pm$2.44 80.57$\pm$3.84 28.57$\pm$1.30 33.07$\pm$0.77 39.91$\pm$0.30 50.27$\pm$1.56
LC No Yes 71.25$\pm$3.17 82.25$\pm$2.11 86.21$\pm$1.02 91.16$\pm$0.97 34.56$\pm$0.69 37.34$\pm$0.69 47.81$\pm$2.00 54.55$\pm$1.26
ERM No No 35.19$\pm$3.49 52.09$\pm$2.88 65.86$\pm$3.59 82.17$\pm$0.74 23.08$\pm$1.25 25.82$\pm$0.33 30.06$\pm$0.71 39.42$\pm$0.64
uLA No No 75.13$\pm$0.78 81.80$\pm$1.41 84.79$\pm$1.10 92.79$\pm$0.85 34.39$\pm$1.14 62.49$\pm$0.74 63.88$\pm$1.07 74.49$\pm$0.58
CCDB No No 83.20$\pm$2.17 87.95$\pm$1.59 91.02$\pm$0.28 96.37$\pm$0.25 55.07$\pm$0.85 63.28$\pm$0.46 67.78$\pm$0.78 74.64$\pm$0.34
Compared Baselines. To demonstrate the superiority of our CCDB in addressing distribution shift in the presence of spurious correlations, we compare it with a diverse set of group robustness techniques. GroupDRO [22] is bias-supervised during both training and validation, serving as a challenging baseline. LFF [24], JTT [25], LC [26], and SELF [2] are pseudo-bias-supervised during training but require bias annotations during validation to achieve optimal performance. Similar to our method, standard empirical risk minimization(ERM), uLA [8], MASKTUNE [30], and XRM [29] are bias-free during both training and validation.
Training setup. To make a fair comparison, we apply a 3-hidden layer MLP for cMNIST, a ResNet18 For cCIFAR10, a ResNet50 for Waterbirds and CelebA, and a BERT for CivilComments following existing methods [8, 25, 29]. Both ResNet18 and ResNet 50 are pretrained on ImageNet-1K, BERT is pretrained on Book Corpus and English Wikipedia. The proposed CCDB method consists of three stages, each optimized independently. The first two training stages span 20 epochs and 1000 iterations across all datasets (except CivilComments with 10 eopchs); the third stage involves 2000 iterations for Waterbirds and CelebA, 5000 iterations for cMNIST and cCIFAR10, and 10 epoches for CivilComments. For Stage1, we use the Adam optimizer with a learning rate of $2\times 10^{-4}$ and default momentum parameters $\beta=(0.9,0.999)$ for all image datasets and AdamW optimizer with a learning rate of $1\times 10^{-5}$ for the NLP dataset; For Stage2, the Adam optimizer with a learning rate of $1\times 10^{-2}$ is adopted across all datasets; For Stage 3, the optimizer, along with its learning rate and other hyperparameters used for each dataset, is detailed in Table II.
On data preprocessing, we apply simple augmentations to each dataset as listed in Table II, following [68]. These augmentations are specifically designed not to induce invariance to variations in the bias attribute. Consequently, the improvements reported in the main paper can be attributed to the proposed CCDB methodology, rather than to the use of bias-nullifying data augmentations.
All the results are averaged over 5 random seeds, with the best-performing model (highest validation accuracy) selected for each seed. We’ll see in Figure 2 that it’s OK for CCDB to use either the standard average accuracy or the worst-class accuracy on a class-balanced validation set for model selection.
TABLE IV: Results on Waterbirds, CelebA and CivilComments. We report average test accuracy(%) and std.dev. over 5 random seeds.
Methods Bias label Waterbirds CelebA CivilComments
Train Val i.i.d. WorstGroup i.i.d. WorstGroup i.i.d. WorstGroup
GroupDRO Yes Yes 93.50 91.40 92.90 88.90 84.2 73.7
LFF No Yes 97.50 75.20 86.00 77.20 68.2 50.3
JTT No Yes 93.60 86.00 88.00 81.10 83.3 64.3
LC No Yes - 90.50$\pm$1.1 - 88.10$\pm$0.8 - 70.30$\pm$1.2
SELF No Yes - 93.00$\pm$0.3 - 83.90$\pm$0.9 - 79.10$\pm$2.1
ERM No No 97.30 72.60 95.60 47.20 81.6 66.7
MASKTUNE No No 93.00$\pm$0.7 86.40$\pm$1.9 91.30$\pm$0.1 78.00$\pm$1.2 - -
uLA No No 91.50$\pm$0.7 86.10$\pm$1.5 93.90$\pm$0.2 86.50$\pm$3.7 - -
XRM No No 90.60 86.10 91.0 88.5 83.5 70.1
CCDB No No 92.59$\pm$0.10 90.48$\pm$0.28 90.08$\pm$0.19 85.27$\pm$0.28 83.60$\pm$0.21 75.00$\pm$0.26

### v3

@@ 6 Related works
6.1 Bias-guided methods
@@
Addressing spurious correlations becomes straightforward when bias annotations are available. By combining both bias and label annotations, one can create a group balanced training process or model selection strategy, to enhance robustness against spurious correlations. For example, Group distributionally robust optimization (GroupDRO) [15, 53, 54, 55] proposes to minimize the worst-group loss based on predefined group assignments. To alleviate the reliance on bias annotations, deep feature reweighting (DFR) [20, 56] proposes retraining only the last layer of the model using a small, group-balanced held-out dataset. MAPLE [21] utilizes a validation set with bias annotations to reweight the training samples, ensuring that the training process focuses on the worst group. Pseudo-bias-supervised methods propose to use the predicted bias information instead of explicit annotations, typically through auxiliary models. Learning from Failure (LFF) [17], logit adjustment technique (LC) [19] and Disentangled Feature Augmentation (DFA) [33] reweight the loss of the unbiased model using a co-trained biased model. Meanwhile, Just-train-twice (JTT) [18] reweights misclassified samples from an initial biased training phase, emphasizing data from the worst-performing groups. Additionally, research in [16, 34, 57] integrates large-scale vision-language models for more precise bias prediction. However, the performance of these methods depends on the prediction quality and still requires bias annotations during validation for optimal performance.
@@
Table 3: An illustration of the datasets used in this paper. In the training data, $\boldsymbol{z}_{\text{s}}$ is highly correlated with the class labels. However, this correlation is either broken or even reversed in the test set.
@@
Table 4: The hyperparameters and augmentations used during the training of our CCDB.
@@
6.2 Bias-agnostic methods
To eliminate the dependence on bias annotations/predictions, researchers have developed bias-agnostic methods under different motivations. Stable learning based methods [35] propose to decorrelate all features to exclude spurious correlations. For example, [31] proposes a sample reweighting approach to decorrelate input variables and demonstrate stable performance across distribution shifts. StableNet [5] further generalizes these techniques to nonlinear feature decorrelation. However, achieving perfect decorrelation theoretically requires a sufficiently large dataset, which is often impractical. To compensate for the imperfectness of sample reweighting under finite-sample settings, Sparse variable independent(SVI) [30] introduces a sparsity constraint to avoid unnecessary decorrelation within core variables. Most of these methods are developed within linear frameworks with predefined features, and their extension to deep models is less effective. This is because the learned features from deep models are often heavily entangled, making decorrelation insufficient to disentangle core and spurious factors [30]. Another line of research utilizes disagreements among diverse models to improve generalization under distribution shifts. For instance, Diversity by disagreement [36] and Diversify and disambiguate [37] propose to maximize the disagreement between multiple predictors to learn a diverse ensemble, which require a completely spuriously correlated training dataset. Selective last-layer finetuning (SELF) [2] proposes reweighting the dataset based on misclassifications or disagreements observed before and after last-layer retraining. However, these methods still need a small set of group annotations for model selection. Several other bias-agnostic methods have demonstrated competitive performance, e.g., General Greedy De-bias Learning [38] leverages both a biased model and a base model to assess the difficulty of training samples, thereby focusing training on the more challenging ones. However, this approach may overestimate the level of bias, necessitating the use of curriculum regularization to strike a better balance between in-distribution and out-of-distribution performance. uLA [7] employs biased prediction to adjust the logits for debias learning. It relies heavily on extensively pretrained self-supervised models, which significantly influence the final performance. XRM [22] constructs twin networks to fully predict bias across the entire training dataset, using a label-flipping mechanism to intentionally steer the model toward biased predictions. MASKTUNE [23], on the other hand, explores new features during a secondary training phase by masking previously identified spurious factors. However, it requires careful selection of the masking threshold to ensure stable performance. The work presented in [58] introduces a novel perspective on the causes of spurious correlations. While it shares certain similarities with our approach, it is grounded in a different framework and primarily aims to offer a comprehensive analysis of existing methods.
Our method also falls within the bias-agnostic framework. In contrast to the aforementioned methods, our approach ($i$) does not decorrelate all features, but instead takes a more straightforward approach by maximizing the conditional entropy of label information given the spurious factors; ($ii$) allows for easy control of sample weighting through distribution distance minimization.
Figure 6: A detailed causal graph.
7 Viewing CCDB from a causal inference perspective
The causal graph in Fig. 1 (left) of the main manuscript is constructed following [32, 42]. A more detailed version is shown in Fig. 6, where blue arrows represent paths whose parameters vary across domains, while black arrows denote domain-invariant paths. Here, $\boldsymbol{u}$ denotes an unobserved confounder (e.g., the domain or environment) that simultaneously affects both $\boldsymbol{z}_{\text{c}}$ and $\boldsymbol{z}_{\text{s}}$, and this causal influence changes across domains. The mechanism by which $\boldsymbol{z}_{\text{c}}$ and $\boldsymbol{z}_{\text{s}}$ jointly generate the observation $\boldsymbol{x}$ is assumed to be invariant across domains. The prediction of $y$ based on $\boldsymbol{z}_{\text{c}}$ is also domain-invariant, whereas the prediction of $y$ based on $\boldsymbol{z}_{\text{s}}$ varies across domains. Treating $\boldsymbol{z}_{\text{c}}$ as the treatment and $y$ as the outcome induces a causal path $\boldsymbol{z}_{\text{c}}\rightarrow y$ as well as a backdoor path $\boldsymbol{z}_{\text{c}}\leftarrow\boldsymbol{u}\rightarrow\boldsymbol{z}_{\text{s}}\rightarrow y$. The spurious correlation between $\boldsymbol{z}_{\text{s}}$ and $y$ arises from the unknown confounder $\boldsymbol{u}$. To safely estimate the causal effect of an image feature $\boldsymbol{z}$ on the label $y$, it is necessary to remove the confounding bias induced by the backdoor path and directly estimate the causal effect of $\boldsymbol{z}_{\text{c}}$ on $y$, which remains invariant across domains, namely,
$p(y|do(\boldsymbol{z}_{\text{c}}))=\mathbb{E}_{p(\boldsymbol{z}_{\text{s}})}p(y|\boldsymbol{z}_{\text{c}},\boldsymbol{z}_{\text{s}})$ (8)
where $do(\boldsymbol{z}_{\text{c}})$ denotes the do-operator from causal inference, which cuts off all incoming paths to $\boldsymbol{z}_{\text{c}}$, thereby removing the spurious correlation between $\boldsymbol{z}_{\text{c}}$ and $\boldsymbol{z}_{\text{s}}$, i.e., $p(\boldsymbol{z}_{\text{s}}|\boldsymbol{z}_{\text{c}})=p(\boldsymbol{z}_{\text{s}})$, a condition that reflects covariate balance. Given that the core factor $\boldsymbol{z}_{\text{c}}$ and the label $y$ have a strong one-to-one correspondence, it follows that $p(\boldsymbol{z}_{\text{s}}|y)=p(\boldsymbol{z}_{\text{s}})$. This condition is equivalent to the optimal solution that minimizes the mutual information term in our CCDB framework.